OpenStudy (anonymous):

FAN AND MEDAL!!!! Choose the equation that represents the graph below: Graph of a line passing through points 0 comma 6 and 9 comma 0 y = -2x - 6 y = negative two thirdsx + 6 y = -negative two thirdsx + 6 y = 2x + 6

OpenStudy (anonymous):

@kl0723

OpenStudy (anonymous):

@pooja195

OpenStudy (misty1212):

HI!!

jimthompson5910 (jim_thompson5910):

You can check each equation. Let's pick choice D at random y = 2x + 6

jimthompson5910 (jim_thompson5910):

plug in (0,6) y = 2x + 6 6 = 2(0) + 6 6 = 6 ... this works ---------------- plug in (9,0) y = 2x + 6 0 = 2(9) + 6 0 = 18 + 6 0 = 24 ... this does not, so the equation cannot be y = 2x + 6

jimthompson5910 (jim_thompson5910):

you'll do the same for each answer choice

jimthompson5910 (jim_thompson5910):

that's one way to do it of course, there is another way

OpenStudy (anonymous):

ok thanks @jim_thompson5910 -2/3x + 6

jimthompson5910 (jim_thompson5910):

let's check that plug in (x,y) = (0,6) y = (-2/3)x + 6 6 = (-2/3)(0) + 6 6 = 0 + 6 6 = 6 ... works ---------------- plug in (x,y) = (9,0) y = (-2/3)x + 6 0 = (-2/3)(9) + 6 0 = -6 + 6 0 = 0 ... works, so the equation is indeed (-2/3)x + 6

OpenStudy (anonymous):

6= 2/3(0) + 6 6= 0 + 6 6=6

OpenStudy (anonymous):

yeah was working on the rest

OpenStudy (anonymous):

wait you plugged in 9 for y but 9 is x

jimthompson5910 (jim_thompson5910):

I replaced y with 0 in the second part

OpenStudy (anonymous):

oh sorry I suck at math

jimthompson5910 (jim_thompson5910):

you're doing great, no worries